# Surface groups among cubulated hyperbolic and one-relator groups

### v1

@@ 3.2 A bound on the widths of $2$-cuts
$\mathrm{Wh}_{X}(Y_{n-1})=\mathrm{Wh}_{X}(Y_{n})\#_{K_{n}}\mathrm{Wh}_{X}(Y^{\prime}_{n})\,,$
@@
We now turn our attention to some CAT(0) geometry of hyperplanes. Consider a 2-cut $Y$ between hyperplanes $H_{1}$ and $H_{2}$, and let $\gamma$ be an $\ell^{2}$ geodesic in between $H_{1}$ and $H_{2}$. In order to prove the desired bound, we need to find long chains of mutually disjoint hyperplanes that cross $\gamma$ at an angle that isn’t too acute. A hyperplane $K$ is said to be $\epsilon$-almost orthogonal to $\gamma$ if $\gamma$ crosses $K$ at an angle of at least $\epsilon$.
@@
Let $d\geq 2$, and let $\gamma$ be an $\ell^{2}$ geodesic in a $d$-dimensional CAT(0) cube complex $\widetilde{X}$. Suppose that $\gamma$ crosses two hyperplanes $K_{1}$ and $K_{2}$ consecutively. Then at least one of $K_{1}$ and $K_{2}$ is $\epsilon_{d}$-almost orthogonal to $\gamma$, where $\epsilon_{d}=\tan^{-1}\left(\frac{1}{2\sqrt{d-1}}\right)$.
@@
The first such hyperplane is by assumption $K_{2}$. Let $\gamma(t_{1})\in K_{2}$ and let $q$ be the point of intersection between $K_{1}$ and $K_{2}$. Then the three points $\gamma(0)$, $\gamma(t_{1})$, $q$ together span a right-angled Euclidean triangle, whence $\gamma$ makes an angle
@@
Let $H_{1}$ and $H_{2}$ be hyperplanes in a $d$-dimensional CAT(0) cube complex $\widetilde{X}$, and let $\gamma$ be an $\ell^{2}$-geodesic between them. For every $n$ there is a $C$ such that, if the length of $\gamma$ is greater than $C$, then there exists a sequence of mutually disjoint hyperplanes
@@
Let $z$ be the closest point on $\gamma$ to $y$, and consider the geodesic triangle with vertices $x$, $y$ and $z$. Let $\bar{\theta}$ be the angle in the Euclidean comparison triangle at the vertex corresponding to $x$, which is at least $\theta$ by [8, Proposition 1.7(4)]. Now, the Euclidean sine rule gives
@@
Let $X$ be a compact, non-positively curved cube complex, and let $R\geq 0$. The set of isomorphism types of abstract bounding hyperplane components $Z$ over $X$ of diameter at most $R$ is finite.
@@
For the converse, let $Y$ be a 2-cut between hyperplanes $H_{1}$ and $H_{2}$. By Lemma 3.6, we may replace $Y$ by its intersection with $\operatorname{Hull}(H_{1}\cup H_{2})$, and so assume that $Y$ is contained in $\operatorname{Hull}(H_{1}\cup H_{2})$. Let $\gamma$ be an $\ell^{2}$ geodesic between $H_{1}$ and $H_{2}$. It is easy to see that $H_{1}\cup\gamma\cup H_{2}$ is $2\delta$-quasiconvex in the $\ell^{2}$ metric and therefore, by Theorem 1.7, $\operatorname{Hull}(H_{1}\cup H_{2})$ is contained in the $R$-neighbourhood of $H_{1}\cup\gamma\cup H_{2}$, for some uniform $R$.
Let $M$ be the number of isomorphism types of oriented abstract hyperplane components over $X$ of diameter at most $2R\sqrt{4d-3}$ (which is finite by Remark 3.10). Let $M^{\prime}$ be such that any geodesic crossing $M^{\prime}$ hyperplanes has an $\ell^{2}$ length of at least $(1+\sqrt{4d-3})R$, and let $N=M+2M^{\prime}$. By Lemma 3.8, if the width of $Y$ is greater than a constant $C$, then we can find a sequence of $M+1$ hyperplanes
@@
From each $K_{i}$ we define an oriented abstract hyperplane component $Z_{i}$ as follows. First, set $Z_{i}=K_{i}\cap Y$; we next need to partition $\pi_{0}(\mathrm{Wh}_{K_{i}}(Z_{i}))$. Cutting $Y$ along $K_{i}$ decomposes it as
@@
In particular, $\mathrm{Wh}_{X}(Y)$ contains a copy of $\mathrm{Wh}_{K_{i}}(Z_{i})$, which is canonically isomorphic to the links of $K_{i}$ in both $\mathrm{Wh}_{X}(Y_{i})$ by Lemma 2.13. The partition on $\pi_{0}(\mathrm{Wh}_{K_{i}}(Z_{i}))$ is now induced by connectivity in $\mathrm{Wh}_{X}(Y)\smallsetminus\{H_{1},H_{2}\}$. Furthermore, each $Z_{i}$ is naturally oriented from $H_{1}$ to $H_{2}$.
@@
Applying Lemma 2.24 twice, $Y^{\prime}$ is a 2-cut between $K_{i}$ and $K_{j}$, while $Z_{i}$ and $Z_{j}$ are the corresponding bounding hyperplane components. Therefore, $Y^{\prime}$ is in fact a periodic 2-cut between $K_{i}$ and $K_{j}$, as required. ∎
@@ 3.3 From $2$-cuts to $k$-cuts
In the previous section, we proved a uniform bound on the widths of 2-cuts. In this section, we will upgrade this to a uniform bound on the widths of $k$-cuts, where the bound is independent of $k$. The key result is Lemma 3.16, which encapsulates the fact that a finite collection of hyperplanes coarsely has the structure of a quasi-tree. I am grateful to Mark Hagen for pointing out that this lemma should be true, and that it would simplify the proof my argument.
The proof of Lemma 3.16 is inspired by the projection-complex machinery of Bestvina–Bromberg–Fujiwara [5]. Consider a collection $\mathbb{V}=\{V_{i}\}$ of convex subcomplexes of a hyperbolic, CAT(0) cube complex $\widetilde{X}$. (In [5] this collection would typically be infinite. For us it will be finite, but of unbounded cardinality.) For each $U\in\mathbb{V}$, we consider the closest-point projection $\pi_{U}:\widetilde{X}\to U$. For each $U$, Bestvina–Bromberg–Fujiwara use this to define a distance function on pairs of elements of $\mathbb{V}\smallsetminus\{U\}$ via
@@
They go on to state projection complex axioms for the distance functions $d^{\pi}_{\bullet}$, with respect to a fixed constant $\theta>0$; see [6, §1.3] for a concise statement of the axioms. The next proposition asserts that, as long as the elements of $\mathbb{V}$ are pairwise sufficiently distant from each other, the projection complex axioms are satisfied.
@@
Let $\widetilde{X}$ be a $\delta$-hyperbolic CAT(0) cube complex and let $\mathbb{V}$ be a finite set of convex subcomplexes of $\widetilde{X}$. There is a constant $C$, depending only on $\widetilde{X}$, such that, as long as the elements of $\mathbb{V}$ are pairwise at distance at least $C$, there is $\theta>0$ such that the following hold for all distinct $U,V,W,Y\in\mathbb{V}$:
@@
    if $d^{\pi}_{U}(V,W)>\theta$ then $d_{V}(U,W)\leq\theta$;
@@
Properties (i) and (ii) are obvious, while property (v) is immediate by hypothesis since $\mathbb{V}$ is assumed to be finite. Property (iii) is known as the contracting property for convex subspaces of $\delta$-hyperbolic spaces, and a standard exercise shows that it holds for any $\theta\geq 8\delta$ as long as $C\geq 2\delta$. Finally, property (iv) is known as the Behrstock inequality. A sketch is given in [44, p. 133]. ∎
@@
Bestvina–Bromberg–Fujiwara–Sisto use the strengthened projection complex axioms to construct a quasitree. Instead, we will use them to find $U\in\mathbb{V}$ onto which the diameter of the projection of all the other elements of $\mathbb{V}$ is bounded.
@@
Let $\widetilde{X}$ be a $\delta$-hyperbolic CAT(0) cube complex and let $\mathbb{V}$ be a finite set of unbounded convex subcomplexes of $\widetilde{X}$. There is a constant $C$, depending only on $\widetilde{X}$, such that, as long as the elements of $\mathbb{V}$ are pairwise at distance at least $C$, there is $U\in\mathbb{Y}$ such that
@@
Take $C$ as in Proposition 3.12. Since $d_{\bullet}$ is within $2\theta$ of $d^{\pi}_{\bullet}$, it suffices to prove that there is $U$ such that $d_{U}(V,W)\leq 11\theta$ for all $V,W\neq U$. We prove this by induction on the cardinality of $\mathbb{V}$.
The base cases when #$\#\mathbb{V}\leq 2$ are trivial. For the inductive step, fix $W_{0}\in\mathbb{V}$ and let $\mathbb{V}^{\prime}:=\mathbb{V}\smallsetminus\{W_{0}\}$. By induction, there is $U^{\prime}\in\mathbb{V}^{\prime}$ such that $d_{U^{\prime}}(V^{\prime},W^{\prime})\leq 11\theta$ for all $V^{\prime},W^{\prime}\in\mathbb{V}^{\prime}\smallsetminus\{U^{\prime}\}$.
If $d_{U^{\prime}}(V^{\prime},W_{0})\leq 11\theta$ for all $V^{\prime}\in\mathbb{V}^{\prime}$ then we are done, taking $U=U^{\prime}$. Suppose, therefore, that $d_{U^{\prime}}(V^{\prime},W_{0})>11\theta$ for some $V^{\prime}\in\mathbb{V}^{\prime}\smallsetminus\{U^{\prime}\}$. In this case, we set $U:=W_{0}$, and we need to show that $d_{U}(V,W)\leq 11\theta$ for any $V,W\in\mathbb{V}^{\prime}$.
Proposition 3.13(iv) implies that
$d_{U^{\prime}}(V,U)=d_{U^{\prime}}(V,W_{0})=d_{U^{\prime}}(V^{\prime},W_{0})>11\theta$
for any $V\neq U^{\prime}$, while $d_{U^{\prime}}(U^{\prime},U)>11\theta$ since $U^{\prime}$ is unbounded. Applying Proposition 3.13(iv) again, it follows that $d_{U}(V,W)=d_{U}(V,U^{\prime})$ for any $W\neq U$. But $d_{U}(V,U^{\prime})\leq 11\theta$ by Remark 3.14, so $d_{U}(V,W)\leq 11\theta$ for any $V,W\neq U$, as required. ∎
@@
Let $\widetilde{X}$ be a $d$-dimensional, $\delta$-hyperbolic CAT(0) cube complex. There is a constant $D$ with the following property. Suppose that $\mathbb{V}$ is any finite, non-empty set of unbounded convex subcomplexes of $\widetilde{X}$. If $d(U,V)>D$ for every distinct $U,V\in\mathbb{V}$, then there is a hyperplane $H$ and $U\in\mathbb{V}$ such that $H$ separates $U$ from every $V\in\mathbb{V}\smallsetminus\{U\}$.
@@
Let $C$ and $\theta$ be as in Lemma 3.15. By Theorem 1.7, there is $r$ such that, for any subset $Z\subseteq\widetilde{X}$ of diameter at most $13\theta$, the diameter of $\operatorname{Hull}(Z)$ is at most $r$.
@@
$\operatorname{diam}\left(\bigcup_{\widehat{V}\in\widehat{\mathbb{V}}\smallsetminus\widehat{U}}\pi_{\widehat{U}}(\widehat{V})\right)\leq 13\theta\,,$
whence
$\operatorname{diam}\operatorname{Hull}\left(\bigcup_{\widehat{V}\in\widehat{\mathbb{V}}\smallsetminus\widehat{U}}\pi_{\widehat{U}}(\widehat{V})\right)\leq r$
by the definition of $r$.
This hull contains a point on the frontier of $\widehat{U}$, so is at distance at least $\sqrt{d}$ from $U$ itself. Therefore, any shortest path from $U$ to the hull crosses some hyperplane $H$. Thus, $H$ separates $U$ from $\pi_{\widehat{U}}(V)$, for every $V\neq U$. It follows that $H$ separates $U$ from $V$, for all $V\neq U$, as required. ∎

### v2

@@ 3.2 A bound on the widths of $2$-cuts
$\mathrm{Wh}_{X}(Y_{n-1})\cong\mathrm{Wh}_{X}(Y_{n})\#_{K_{n}}\mathrm{Wh}_{X}(Y^{\prime}_{n})\,,$
@@
We now turn our attention to some CAT(0) geometry of hyperplanes. Consider a 2-cut $Y$ between hyperplanes $H_{1}$ and $H_{2}$, and let $\gamma$ be an $\ell^{2}$-geodesic between $H_{1}$ and $H_{2}$. In order to prove the desired bound, we need to find long chains of mutually disjoint hyperplanes that cross $\gamma$ at an angle that isn’t too acute. A hyperplane $K$ is said to be $\epsilon$-almost orthogonal to $\gamma$ if $\gamma$ crosses $K$ at an angle of at least $\epsilon$.
@@
Let $d\geq 2$, and let $\gamma$ be an $\ell^{2}$-geodesic in a $d$-dimensional CAT(0) cube complex $\widetilde{X}$. Suppose that $\gamma$ crosses two hyperplanes $K_{1}$ and $K_{2}$ consecutively. Then at least one of $K_{1}$ and $K_{2}$ is $\epsilon_{d}$-almost orthogonal to $\gamma$, where $\epsilon_{d}=\tan^{-1}\left(\frac{1}{2\sqrt{d-1}}\right)$.
@@
The first such hyperplane is by assumption $K_{2}$. Let $\gamma(t_{1})\in K_{2}$ and let $q$ be the orthogonal projection of $\gamma(0)$ to $K_{2}$. Then the three points $\gamma(0)$, $\gamma(t_{1})$, $q$ together span a right-angled Euclidean triangle, whence $\gamma$ makes an angle
@@
Let $H_{1}$ and $H_{2}$ be hyperplanes in a $d$-dimensional CAT(0) cube complex $\widetilde{X}$, and let $\gamma$ be a shortest $\ell^{2}$-geodesic between them. For every $n$ there is a $C$ such that, if the length of $\gamma$ is greater than $C$, then there exists a sequence of mutually disjoint hyperplanes
@@
Let $z$ be the closest point on $\gamma$ to $y$, and consider the geodesic triangle with vertices $x$, $y$ and $z$. Let $\bar{\theta}$ be the angle in the Euclidean comparison triangle at the vertex corresponding to $x$, which is at least $\theta$ by [10, Proposition 1.7(4)]. Now, the Euclidean sine rule gives
@@
Let $X$ be a compact, non-positively curved cube complex, and let $R\geq 0$. The set of isomorphism types of abstract bounding hyperplane components $Z$ over $X$ of diameter at most $R$ is finite. Indeed, the bound on diameter implies that there are only finitely many possible subcomplexes $Z$ of hyperplanes $H$ of $\widetilde{X}$, up to the action of $\pi_{1}(X)$; since $\mathrm{Wh}_{H}(Z)$ is compact for each $Z$, the number of possible cohomology classes in $\tilde{H}^{0}(\mathrm{Wh}_{H}(Z);\mathbb{Z}/2\mathbb{Z})$ is also finite.
@@
For the converse, let $Y$ be a 2-cut between hyperplanes $H_{1}$ and $H_{2}$. By Lemma 3.7, we may replace $Y$ by its intersection with $\operatorname{Hull}(H_{1}\cup H_{2})$, and so assume that $Y$ is contained in $\operatorname{Hull}(H_{1}\cup H_{2})$. Let $\gamma$ be an $\ell^{2}$-geodesic between $H_{1}$ and $H_{2}$. It is easy to see that $H_{1}\cup\gamma\cup H_{2}$ is $2\delta$-quasiconvex in the $\ell^{2}$ metric and therefore, by Theorem 1.7, $\operatorname{Hull}(H_{1}\cup H_{2})$ is contained in the $R$-neighbourhood of $H_{1}\cup\gamma\cup H_{2}$, for some uniform $R$.
Let $M$ be the number of isomorphism types of oriented abstract hyperplane components over $X$ of diameter at most $2R\sqrt{4d-3}$ (which is finite by Remark 3.11). Let $M^{\prime}$ be such that any geodesic crossing $M^{\prime}$ hyperplanes has an $\ell^{2}$-length of at least $(1+\sqrt{4d-3})R$, and let $N=M+2M^{\prime}$. By Lemma 3.9, if the width of $Y$ is greater than a constant $C$, then we can find a sequence of $M+1$ hyperplanes
@@
From each $K_{i}$ we define an oriented abstract hyperplane component $Z_{i}$ as follows. After setting $Z_{i}=K_{i}\cap Y$, we need to define a non-trivial cohomology class $c_{Z_{i}}\in\tilde{H}^{0}(\mathrm{Wh}_{K_{i}}(Z_{i}))$ where, as usual, we consider cohomology with coefficients in $\mathbb{Z}/2\mathbb{Z}$. Cutting $Y$ along $K_{i}$ decomposes it as
@@
In particular, $\mathrm{Wh}_{X}(Y)_{\{H_{1},H_{2}\}}$ contains a copy of $L_{i}=\mathrm{Wh}_{K_{i}}(Z_{i})$, which is canonically isomorphic to the links of $K_{i}$ in both $\mathrm{Wh}_{X}(Y_{1})$ and $\mathrm{Wh}_{X}(Y_{2})$ by Lemma 2.13. Since $Y$ is a 2-cut, there is a non-trivial class $c_{Y}$ in the reduced cohomology group $\tilde{H}^{0}(\mathrm{Wh}_{X}(Y)_{\{H_{1},H_{2}\}})$. The cohomology class $c_{Z_{i}}$ is then defined as the pullback of $c_{Y}$ along the inclusion of $L_{i}$ into $\mathrm{Wh}_{X}(Y)_{\{H_{1},H_{2}\}}$. We claim that $c_{Z_{i}}\neq 0$.
Indeed, Mayer–Vietoris applied to the decomposition
$\mathrm{Wh}_{X}(Y)_{\{H_{1},H_{2}\}}=\mathrm{Wh}_{X}(Y_{1})_{\{H_{1},K_{i}\}}\cup_{L_{i}}\mathrm{Wh}_{X}(Y_{2})_{\{K_{i},H_{2}\}}$
shows that the pullback $b$ of $c_{Y}$ to $\tilde{H}^{0}(\mathrm{Wh}_{X}(Y_{1})_{\{H_{1},K_{i}\}})$ (without loss of generality) is non-zero. But $c_{Z_{i}}$ is the pullback of $b$ to $\tilde{H}^{0}(L_{i})$, so if $c_{Z_{i}}=0$ then, by Mayer–Vietoris applied to the decomposition
$\mathrm{Wh}_{X}(Y_{1})_{H_{1}}=\mathrm{Wh}_{X}(Y_{1})_{\{H_{1},K_{i}\}}\cup_{L_{i}}C(L_{i})\,,$
we see that $b$ is the image of a non-trivial cohomology class in $\tilde{H}^{0}(\mathrm{Wh}_{X}(Y_{1})_{H_{1}})$, contradicting the hypothesis that $X$ has no 1-cuts. Hence, $c_{Z_{i}}\neq 0$ as claimed.
@@
Applying Lemma 2.24 twice, $Y^{\prime}$ is a 2-cut between $K_{i}$ and $K_{j}$, while $Z_{i}$ and $Z_{j}$ are the corresponding bounding hyperplane components.
Finally, note that the inclusions of $L_{i}=\mathrm{Wh}_{K_{i}}(Z_{i})$ and $L_{j}=\mathrm{Wh}_{K_{j}}(Z_{j})$ into $\mathrm{Wh}_{X}(Y)_{\{H_{1},H_{2}\}}$ factor through the inclusion of $\mathrm{Wh}_{X}(Y^{\prime})_{\{K_{i},K_{j}\}}$. Therefore, setting $c_{Y^{\prime}}$ to be the pullback of $c_{Y}$ to $\mathrm{Wh}_{X}(Y^{\prime})_{\{K_{i},K_{j}\}}$, we see that $c_{Y^{\prime}}$ pulls back to both $c_{Z_{i}}$ and $c_{Z_{j}}$. This completes the proof that $Y^{\prime}$ is a periodic 2-cut. ∎
@@ 3.3 From $2$-cuts to $k$-cuts
In the previous section, we proved a uniform bound on the widths of 2-cuts. In this section, we will upgrade this to a uniform bound on the widths of $k$-cuts, where the bound is independent of $k$. The key result is Lemma 3.20, which encapsulates the fact that a finite collection of hyperplanes coarsely has the structure of a quasi-tree. I am grateful to Mark Hagen for pointing out that this lemma should be true, and that it would simplify the proof of my argument.
The proof of Lemma 3.20 is inspired by the projection-complex machinery of Bestvina–Bromberg–Fujiwara [6] [^5] Anthony Genevois has communicated to me a shorter, direct proof of the lemma that avoids the Bestvina–Bromberg–Fujiwara machinery [25].. Consider a collection $\mathbb{V}=\{V_{i}\}$ of convex subcomplexes of a hyperbolic, CAT(0) cube complex $\widetilde{X}$. (In [6] this collection would typically be infinite. For us it will be finite, but of unbounded cardinality.) For each $U\in\mathbb{V}$, we consider the closest-point projection $\pi_{U}:\widetilde{X}\to U$. For each $U$, Bestvina–Bromberg–Fujiwara use this to define a distance function on pairs of elements of $\mathbb{V}\smallsetminus\{U\}$ via
@@
They go on to state projection complex axioms for the distance functions $d^{\pi}_{\bullet}$, with respect to a fixed constant $\theta>0$; see [7, §1.3] for a concise statement of the axioms. The next proposition asserts that, as long as the elements of $\mathbb{V}$ are pairwise sufficiently distant from each other, the projection complex axioms are satisfied.
@@
Let $\widetilde{X}$ be a $\delta$-hyperbolic CAT(0) cube complex and let $\mathbb{V}$ be a finite set of convex subcomplexes of $\widetilde{X}$. There are constants $C$ and $\theta>0$, depending only on $\widetilde{X}$, such that, as long as the elements of $\mathbb{V}$ are pairwise at distance at least $C$, the following hold for all distinct $U,V,W,Y\in\mathbb{V}$:
@@
    if $d^{\pi}_{U}(V,W)>\theta$ then $d^{\pi}_{V}(U,W)\leq\theta$;
@@
Properties (i) and (ii) are obvious, while property (v) is immediate by hypothesis since $\mathbb{V}$ is assumed to be finite. Property (iii) is known as the contracting property for quasi-convex subspaces of $\delta$-hyperbolic spaces, and a standard exercise shows that it holds for any $\theta\geq 8\delta$ as long as $C\geq 2\delta$. Finally, property (iv) is known as the Behrstock inequality. A sketch is given in [57, p. 133]. ∎
@@
Bestvina–Bromberg–Fujiwara–Sisto use the strengthened projection complex axioms to construct an infinite quasitree $\mathcal{P}_{K}(\mathbb{V})$. In our situation, $\mathcal{P}_{K}(\mathbb{V})$ will be finite, and we will use it to find $U\in\mathbb{V}$ onto which the diameter of the projection of all the other elements of $\mathbb{V}$ is bounded.
Definition 3.16 (Projection complex).
For any $K>0$, the projection complex $\mathcal{P}_{K}(\mathbb{V})$ is a graph with vertex set $\mathbb{V}$. Distinct vertices $U,V$ are joined by an edge if $d_{W}(U,V)\leq K$ for all $W\neq U,V$.
For large enough $K$, they go on to define a standard path in $\mathcal{P}_{K}(\mathbb{V})$ between any $U,V\in\mathbb{V}$; we will let $[U,V]$ denote this standard path. Similarly, we will let $(U,V)$ denote $[U,V]\smallsetminus\{U,V\}$, etc. (In [7], the set of vertices in $(U,V)$ is denoted by $\mathbf{Y}_{K}(U,V)$.) We summarise the properties of standard paths in the following proposition.
Proposition 3.17 (Properties of standard paths).
Suppose $K\geq 33\theta$. Let $U,V\in\mathbb{V}$. The standard path $[U,V]$ is an embedded interval in $\mathcal{P}_{K}(\mathbb{V})$ between $U$ and $V$. Furthermore, the following hold.
  1. (i)
    If $U^{\prime},V^{\prime}\in[U,V]$ then $[U^{\prime},V^{\prime}]\subseteq[U,V]$.
  2. (ii)
    If $n$ is the length of the standard path $[U,V]$ then the shortest path in $\mathcal{P}_{K}(\mathbb{V})$ from $U$ to $V$ has length at least $\lfloor\frac{n}{2}\rfloor+1$.
  3. (iii)
    If $W\in(U,V]$ then $d_{U}(V,W)\leq K$.
Proof.
The proofs of [7, §2–3] use a parameter $\theta$ that satisfies the strong projection axioms. In our context, the strong projection axioms are provided by Proposition 3.14, so we must replace $\theta$ by $11\theta$.
The vertices of $(U,V)$ consist of those $W\in\mathbb{V}$ such that $d_{W}(U,V)>K$. Setting $W_{1}<W_{2}$ if $d_{W_{1}}(U,W_{2})>K$ defines a total order on the vertices of $(U,V)$ by [7, Proposition 2.3], to which we may add $U$ as the least element and $V$ as the greatest element; [7, Lemma 3.1] asserts that these are indeed the vertices of an embedded path.
Item (i) now follows immediately from [7, Corollary 2.5] and item (ii) is [7, Lemma 3.7]. For item (iii), there is nothing to prove if $W=V$; otherwise $d_{W}(U,V)>K>11\theta$ so $d_{U}(V,W)\leq 11\theta<K$ by Remark 3.15. ∎
Bestvina–Bromberg–Fujiwara–Sisto also provide a very precise description of triangles of standard paths.
Lemma 3.18 (Triangles of standard paths).
Suppose $K\geq 33\theta$. Let $U,V,W\in\mathbb{V}$. The complement
$[U,V]\smallsetminus([U,W]\cup[W,V])$
in $\mathcal{P}_{K}(\mathbb{V})$ is connected, and contains at most two vertices.
Proof.
This is [7, Lemma 3.6]. ∎
With these facts in hand, we can prove our desired bound on diameters of projections.
@@
Let $\widetilde{X}$ be a $\delta$-hyperbolic CAT(0) cube complex and let $\mathbb{V}$ be a finite set of unbounded convex subcomplexes of $\widetilde{X}$ with at least two elements. There is a constant $C$, depending only on $\widetilde{X}$, such that, as long as the elements of $\mathbb{V}$ are pairwise at distance at least $C$, there is $U\in\mathbb{Y}$ such that
@@
Let $C$ and $\theta$ be as in Proposition 3.13, and consider the projection complex $\mathcal{P}_{33\theta}(\mathbb{V})$, defined using the distance functions $d_{\bullet}$ from Proposition 3.14. Since $\mathbb{V}$ is finite, there are finitely many standard paths; fix $[U,V]$ of maximal length. Let $V^{\prime}$ be the vertex of $[U,V]$ adjacent to $U$, and let $W^{\prime}$ be any other vertex of $\mathcal{P}_{33\theta}(\mathbb{V})$ adjacent to $U$.
First, note that $U$ is not contained in the standard path from $W^{\prime}$ to $V$. Indeed, if $U$ were contained in $[W^{\prime},V]$ then, by Proposition 3.17(i), $[U,V]$ would be a proper sub-path of $[W^{\prime},V]$, contradicting the choice of $U$. Note also that the standard path $[U,W^{\prime}]$ is a single edge, by Proposition 3.17(ii).
Next, consider the triangle of standard paths $[U,W^{\prime}]\cup[W^{\prime},V]\cup[V,U]$. Combining the facts from the previous paragraph with Lemma 3.18, we see that $V^{\prime}$ is joined to $W^{\prime}$ by an edge-path of length at most 5, avoiding $U$. By the definition of $\mathcal{P}_{33\theta}(\mathbb{V})$ and induction on the length of this edge-path, it follows that $d_{U}(V^{\prime},W^{\prime})\leq 5\times 33\theta=165\theta$.
Now consider an arbitrary $W\neq U$ in $\mathbb{V}$ and let $W^{\prime}$ be the vertex of $[W,U]$ adjacent to $U$. Combining the estimate from the previous paragraph with Proposition 3.17(iii) gives
$d_{U}(V^{\prime},W)\leq d_{U}(V^{\prime},W^{\prime})+d_{U}(W^{\prime},W)\leq 165\theta+33\theta=198\theta\,.$
It follows from Proposition 3.14 that $d^{\pi}_{U}(V^{\prime},W)\leq 200\theta$, and so $d^{\pi}_{U}(W_{1},W_{2})\leq 400\theta$ for any $W_{1},W_{2}\neq U$ in $\mathbb{V}$, by the triangle inequality. This is the required bound, by the definition of $d^{\pi}_{\bullet}$. ∎
@@
Let $\widetilde{X}$ be a $d$-dimensional, $\delta$-hyperbolic CAT(0) cube complex. There is a constant $D$ with the following property. Suppose that $\mathbb{V}$ is any finite, non-empty set of unbounded convex subcomplexes of $\widetilde{X}$ and that $d(U,V)>D$ for every distinct $U,V\in\mathbb{V}$. Then, there exists a hyperplane $H$ and $U\in\mathbb{V}$ such that $H$ separates $U$ from every $V\in\mathbb{V}\smallsetminus\{U\}$.
@@
Let $C$ and $\theta$ be as in Lemma 3.19. By Theorem 1.7, there is $r$ such that, for any subset $Z\subseteq\widetilde{X}$ of diameter at most $400\theta$, the diameter of $\operatorname{Hull}(Z)$ is at most $r$.
@@
$\operatorname{diam}\left(\bigcup_{\widehat{V}\in\widehat{\mathbb{V}}\smallsetminus\widehat{U}}\pi_{\widehat{U}}(\widehat{V})\right)\leq 400\theta\,.$
Let $W\subseteq\widehat{U}$ be the convex hull of $\bigcup_{\widehat{V}\in\widehat{\mathbb{V}}\smallsetminus\widehat{U}}\pi_{\widehat{U}}(\widehat{V})$, so $\operatorname{diam}W\leq r$.
Since $W$ contains a point on the frontier of $\widehat{U}$, it follows from the bound on the diameter of $W$ and the definition of $\widehat{U}$ that $d(U,W)\geq\sqrt{d}$. In particular, by Lemma 1.6 there is a hyperplane $H$ that separates $U$ from $W$; note that $H$ intersects $\widehat{U}$ non-trivially by construction.
Finally, consider any $V\neq U$ in $\mathbb{V}$. It remains to prove that $H$ separates $U$ from $V$. Since $H$ separates $U$ from $W$, it suffices to prove that $H$ is disjoint from $V$ and does not separate $W$ from $V$. Because hyperplanes are convex and $H$ intersects $\widehat{U}$, if $H$ also intersected $V$ then it would intersect $W$, a contradiction. So $H$ is disjoint from $V$. If $H$ separated $W$ from $V$ then a shortest geodesic $\gamma$ from $V$ to $W$ would cross $H$ by Lemma 1.6; but $\gamma$ is also a shortest geodesic from $V$ to $\widehat{U}$, so $H$ would separate $V$ from $\widehat{U}$, contradicting the fact that $H$ intersects $\widehat{U}$. ∎